Costco Wholesale Corporation operated 715 stores as of August 28, 2016. The following data were taken from the company’s annual report. All dollar amounts are in millions.

Fiscal Years Ending

August 28, 2016

August 30, 2015

Net sales

  $116,073

$113,666

Cost of goods sold

102,901

101,065

Net income

2,350

2,377

Merchandise inventory

8,969

   8,908

Page 312Required

  1. Compute Costco’s inventory turnover ratio for 2016 and 2015.
  2. Compute Costco’s average days to sell inventory for 2016 and 2015.
  3. Based on your computations in Requirements a and b, did Costco’s inventory management get better or worse from 2015 to 2016?
Answer #1

Inventory Turnover = Cost of Goods Sold /Average inventory

2016 = 102,901 / [(8969+8908)/2] = 11.51 times

2015 = 101,065/8908 = 11.35 times

Average days to sell = 365/ Inventory Turnover

2016 = 365/11.51 = 31.7 days

2015 = 365/11.35 = 32.12 days

Inventory management got better because inventory Turnover increased

NOTE : I've rounded some numbers. Comment if you face any issues